OpenStudy (anonymous):

write the equation of the line in slope-inercept form that passes through the points (-2,6) and (-1,4) PLEASEE HELP, Will give medal.

OpenStudy (ivettef365):

First you need to find the slope, do you know how to?

OpenStudy (anonymous):

no

OpenStudy (ivettef365):

use formula: y2 - y1 ------- x2 - x1 you have points (-2,6) -> (x1,y1) and (-1,4) -> (x2,y2)

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

wait im confused is it suppose to be written -1-6 for the first one?

OpenStudy (anonymous):

for the top

OpenStudy (anonymous):

is it -1-6/4--2 ?

OpenStudy (anonymous):

is the slope -7/6 ?

OpenStudy (amistre64):

i tend to misplace the values ... so i do a different setup and move the points so that one of the is at the origin; then i can read the slope as y/x off the other point (-2,6) and (-1,4) +2-6 +2-6 --------------- 0 0 1 -2 , its not going to be -7/6

OpenStudy (anonymous):

OK, im sorry but that just confused me more,

OpenStudy (amistre64):

how so?

OpenStudy (anonymous):

the way you set it up, looks as if i have to add everything

OpenStudy (amistre64):

there is addition involved; but the setup is just organizing it all

OpenStudy (amistre64):

im subtracting the parts of one point, from the parts of the other is all

OpenStudy (amistre64):

i know that in order to zero out (-2,6), i have to add 2 and subtract 6: (-2+2,6-6) = (0,0) ; what you do to one point, do to the other (-1+2,4-6) = (1,-2) .... now slope is defined as y/x

OpenStudy (anonymous):

oooh, ok i get you, no one ever taught me that way

OpenStudy (amistre64):

yeah, its just the mechanics behind the formula :)

OpenStudy (anonymous):

ok, lets not talk in mathematical terms, because then your going to loose me again LOL

OpenStudy (amistre64):

fine :) can you define the slope using the point (1,-2) now?

OpenStudy (anonymous):

nope, but i understood what you did before.

OpenStudy (amistre64):

when given a point (x,y), we cen define the slope from the origin as: y/x

OpenStudy (anonymous):

so am i suppose to -1 and add 2

OpenStudy (amistre64):

(1,-2) has the slope: -2/1, or simply, -2

OpenStudy (anonymous):

oh ok, so you just switched x,y to y,x

OpenStudy (anonymous):

?

OpenStudy (amistre64):

visually, yes ... and divided them

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

so you agree that the slope of (1,-2) from the origin is: -2/1, or just -2 right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

now we just have to form the proper equation that they are asking for ... slope intercept

OpenStudy (anonymous):

yea, you need to like break that down in nano details because i dont know how to do that either

OpenStudy (amistre64):

with the information we have, we can form the "\(point\)-\(slope\)" form and work it into the "\(slope\)-\(intercept\)"

OpenStudy (anonymous):

umm mhm

OpenStudy (amistre64):

which one of the given points is your favorite? since it doesnt matter which one we need to use ....

OpenStudy (anonymous):

im guessing Point slope, i dont know what either of them do by the way

OpenStudy (amistre64):

we are given the points (-2,6) and (-1,4); we can use either point, and the slope that we found, to form what is called the: "\(point-slope\)" notice that this from uses a POINT and the SLOPE, hence the name ....

OpenStudy (anonymous):

so we subtracting the point and slope? The slope we just got and the pint i still dont no

OpenStudy (anonymous):

Point*

OpenStudy (amistre64):

thats a hypen (-) , not a subtraction .... its just the name of the format that we can build with the information we have: the format uses as point (a,b), and a slope (m) to build the equation: \(y=m(x-a)+b\)

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

lets use one of the points given: (-2,6); and the slope that we found (-2) and build this format: \[y=m(x-a)+b~~\to\to\to~~y=-2(x-(-2))+6\] or simply\[y=-2(x+2)+6\] we are almost done ...

OpenStudy (anonymous):

so all i know right now is M is supposed to be changed to -2

OpenStudy (anonymous):

ok i got you

OpenStudy (amistre64):

good, m changes to -2, and (a,b) is either point ... i just used the closest one :) now we want to work this int the \(slope-intercept\) form which has the format: \(y=mx+c\)

OpenStudy (anonymous):

isnt it supposed to be b instead of c?

OpenStudy (anonymous):

y=mx+b ??

OpenStudy (amistre64):

since i already used "b" while naming the point, i really didnt want to use it to define the constant ... so i renamed the constant as "c" instead. But yes, the textbooks in america tend to use y=mx+b, while i believe the brits use: y = mx+c .... the numbers really dont care what they are named :) so, let just distribute the slope thru the paranthesis; and add the constants ... \[y=-2(x+2)+6\] \[y=-2x+2(-2)+6\] \[y=-2x-4+6\] \[y=-2x+2\]

OpenStudy (anonymous):

oooh ok, i get you

OpenStudy (amistre64):

we could just have easily used some greek letters :) \[\Large \gamma=\mu\chi+\rho\]

OpenStudy (anonymous):

Yes but then i would have been totally lost lol

OpenStudy (amistre64):

lol, when the teacher asked what to name the variable, I would ask it to be in celtic runes

OpenStudy (amistre64):

the slope-intercept form is defined by its use of the slope (-2), and the "y-intercept" which in our case turns out to be (+2)

OpenStudy (anonymous):

ok :) i get you, but you cant be using your big words like Celtic runes, because then i become stuck on dumb llike what?

OpenStudy (amistre64):

ill try :) good luck